# On Random Fields Associated with Analytic Wavelet Transform

### v1

@@ 1. Introduction
In this work, we study the statistical properties of $|W_{Y}|$ and $\arg(W_{Y})$ when $f$ satisfies appropriate regularity conditions and $\Phi$ is a stationary Gaussian process [18] with mild conditions. Our contributions are multifaceted. First, we derive the joint and marginal probability density functions of $|W_{Y}(t,s)|$ and $\arg(W_{Y}(t,s))$ for any $(t,s)\in\mathbb{R}\times(0,\infty)$. These results yield concentration inequalities quantifying how rapidly the magnitudes and phases of the AWT of $Y$ converge to those of $f$ as the SNR increases, and how the convergence depends on time and scale. Second, we analyze the covariance structure of $|W_{Y}|$ and $\arg(W_{Y})$. In the null case $f=0$, we derive explicit expressions for the covariances of $|W_{Y}|$ and $\arg(W_{Y})$. Using Kolmogorov’s consistency theorem, we show that local covariance structures across discrete time-scale points extend to valid global processes. This provides a theoretical foundation for empirical covariance estimation and supports statistical procedures such as ridge detection and uncertainty quantification. Furthermore, we show the independence of $|W_{Y}|$ and $\arg(W_{Y})$ at different time-scale points. Finally, we derive an SNR-dependent upper bound on the probability of misidentifying the time-scale ridge of $f$ using $|W_{Y}|$, extending our previous ridge analysis in [19], and study the regularity of scalogram contours, showing that almost all contours are differentiable.
AWT $W_{Y}$ of a Gaussian process $Y$ Section 2 distribution of $W_{Y}$ Section 3.2 phase $\arg(W_{Y})$ Section 3.1 magnitude $|W_{Y}|$ Section 3.3 covariance j.p.d.f. (Prop. 1 Cor. 1) contour (Prop. 4) j.p.d.f. (Prop. 5 Cor. 2) stability analysis (Prop. 2-3) stability analysis (Cor. 3) $|W_{Y}|$ (Prop. 6-7 Cor. 4) $\arg(W_{Y})$ (Cor. 5-6) $|W_{Y}|$ vs $\arg(W_{Y})$ (Cor. 7)
@@
The rest of the paper is organized as follows. In Section 2, we present some necessary background, including a review of the continuous wavelet transform and the spectral representation of stationary Gaussian processes. As illustrated in Figure 1, our main findings are presented in Section 3, which is further divided into three subsections addressing: (1) the distributional properties of the magnitude of AWT, (2) the distributional properties of the phase of AWT, and (3) the covariance structure of the AWT’s magnitude and phase. Section 4 concludes the paper, and Section 5 contains all proofs.
@@ 2. Preliminary
$p\left(\mathbf{w};\mathbf{W}_{Y}\right)$
(8) $=$ $\frac{1}{\pi^{n}\textup{det}(\Gamma)}\textup{exp}\left(-\left(\mathbf{w}-\mathbf{W}_{f}\right)^{*}\Gamma^{-1}\left(\mathbf{w}-\mathbf{W}_{f}\right)\right),$
@@ 3.1. Marginal distributions of AWT magnitude
$p\left(r_{1:n};|\mathbf{W}_{Y}|\right)$
(18) $=$ $\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)\int_{0}^{2\pi}\hskip-4.26773pt\int_{0}^{2\pi}\hskip-5.69046pt\cdots\hskip-4.26773pt\int_{0}^{2\pi}\hskip-2.84544ptp\left(\mathbf{w};\mathbf{W}_{Y}\right)d\theta_{1}d\theta_{2}\cdots d\theta_{n}$
@@
(20) $p\left(r_{1:n};|\mathbf{W}_{Y}|\right)$
$=$ $\frac{1}{\pi^{n}\textup{det}(\Gamma)}\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)J\left(\textup{diag}\left(r_{1:n}\right)\Gamma^{-1}\textup{diag}\left(r_{1:n}\right)\right),$
@@
(22) $p\left(r_{1:2};|\mathbf{W}_{\Phi}|\right)$
$=$ $\frac{4r_{1}r_{2}}{\textup{det}(\Gamma)}\exp\left(-\left(\frac{r_{1}^{2}}{\Gamma_{11}}+\frac{r_{2}^{2}}{\Gamma_{22}}\right)\frac{\Gamma_{11}\Gamma_{22}}{\textup{det}(\Gamma)}\right)I_{0}\left(\frac{2|\Gamma_{12}|r_{1}r_{2}}{\textup{det}(\Gamma)}\right),$
@@
$p\left(r;\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}\right)$
(23) $=$ $2q(t,s)re^{-q(t,s)(1+r^{2})}I_{0}\left(2q(t,s)r\right)$
@@
$p\left(r;\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}\right)$
$=$ $2q(t,s)re^{-q(t,s)(1+r^{2})}\frac{e^{2q(t,s)r}}{\sqrt{4\pi q(t,s)r}}\left(1+\mathcal{O}\left(\frac{1}{2q(t,s)r}\right)\right)$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}-1\right|>\varepsilon\right)$
$\leq$ $\textup{exp}\left(-\frac{\varepsilon^{2}}{16}q(t,s)\right)+\textup{exp}\left(-\frac{\varepsilon}{2}q(t,s)\right).$
@@ 3.2. Marginal distributions of AWT phase
(27) $p\left(\theta_{1:n};\Theta_{Y}\right)$
$=$ $\int_{0}^{\infty}\hskip-4.26773pt\int_{0}^{\infty}\hskip-5.69046pt\cdots\hskip-4.26773pt\int_{0}^{\infty}\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)\hskip-2.84544ptp\left(\mathbf{w};\mathbf{W}_{Y}\right)dr_{1}dr_{2}\cdots dr_{n}$
@@
(28) $p(\theta;\Theta_{Y}(t,s))=\frac{1}{2\pi}e^{-q}$
$+\frac{\sqrt{q}}{2\sqrt{\pi}}\cos(\theta-\theta_{f})\textup{exp}\left\{\hskip-1.42271pt-q\left[1-\cos^{2}(\theta-\theta_{f})\right]\hskip-1.42271pt\right\}\hskip-2.84544ptB_{q}(\theta-\theta_{f}),$
@@
(30) $p\left(\theta_{1:2};\Theta_{\Phi}\right)=\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)}{D(\theta_{2}-\theta_{1})}$
$+\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)C(\theta_{2}-\theta_{1})}{D(\theta_{2}-\theta_{1})^{3/2}}\Big{[}\frac{\pi}{2}+\tan^{-1}\Big{(}\frac{C(\theta_{2}-\theta_{1})}{\sqrt{D(\theta_{2}-\theta_{1})}}\Big{)}\Big{]},$
@@
$\mathbb{P}\left(d\left(\Theta_{Y}(t,s),\arg(W_{f}(t,s))\right)>\varepsilon\right)\leq\frac{\varepsilon}{\pi}e^{-q}$
$+$ $\frac{2\varepsilon\sqrt{q}}{\sqrt{\pi}}\cos(\varepsilon)\textup{exp}\left\{-q\left[1-\cos^{2}(\varepsilon)\right]\right\},$
@@
$d\left(\Theta_{Y}(t,s),\arg(W_{f}(t,s))\right)$
$=$ $1-\left|\cos\big{(}\Theta_{Y}(t,s)-\arg(W_{f}(t,s))\big{)}\right|.$
@@ 3.3. Covariance of AWT’s magnitudes and phases
$=$ $4\underset{\Box,\triangle\in\{\Re,\Im\}}{\sum}\Box(W_{f}(t_{1},s_{1}))\triangle(W_{f}(t_{2},s_{2}))$
$\times\sqrt{s_{1}s_{2}}\int_{\mathbb{R}}e^{i(t_{1}-t_{2})\lambda}\overline{\widehat{\Box(\psi)}}(s_{1}\lambda)\widehat{\triangle(\psi)}(s_{2}\lambda)F(d\lambda)$
@@
$\textup{Cov}\left(|W_{Y}(t_{1},s_{1})|,|W_{Y}(t_{2},s_{2})|\right)$
$=$ $\int_{\mathbb{C}}|w|p\left(w;W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)dw$
(32) $-\overset{2}{\underset{\ell=1}{\prod}}\int_{\mathbb{C}}|w|p\left(w;W_{Y}(t_{\ell},s_{\ell})\right)dw,$
@@
(33) $p\left(w;W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)=\int_{0}^{\infty}\int_{0}^{2\pi}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}$
(38) $\times\textup{exp}\left(-\left(\left[\begin{array}[]{cc}re^{i\theta}\\ wr^{-1}e^{-i\theta}\end{array}\right]-\left[\begin{array}[]{cc}W_{f}(t_{1},s_{1})\\ W_{f}(t_{2},s_{2})\end{array}\right]\right)^{*}\right$.
@@
$\textup{Corr}\left(\left|W_{\Phi}(t_{1},s_{1})\right|,\left|W_{\Phi}(t_{2},s_{2})\right|\right)$
(45) $\sim$ $\left(\frac{16}{\pi}-4\right)^{-1}\frac{|\Gamma_{12}|^{2}}{\Gamma_{11}\Gamma_{22}}.$
@@
    $\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}=$ $2^{2\alpha+1}\frac{\left(s_{1}s_{2}\right)^{\alpha+\frac{1}{2}}}{(s_{1}+s_{2})^{2\alpha+1}}$
    $=$ $\left(\frac{2^{\Delta/2}+2^{-\Delta/2}}{2}\right)^{-(2\alpha+1)}.$
@@
    $\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}=$ $2^{2\alpha+1}\frac{s^{2\alpha+1}}{\left|2s-i(t_{1}-t_{2})\right|^{2\alpha+1}}$
    $=$ $\left[1+\left(\frac{t_{1}-t_{2}}{2s}\right)^{2}\right]^{-(\alpha+1/2)}.$
@@
$=$ $\int_{\mathbb{C}}\frac{w}{|w|}p\left(w;W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)dw.$

### v2

@@ 1. Introduction
[…]
AWT $W_{Y}$ of a Gaussian process $Y$ Section 2 distribution of $W_{Y}$ Section 3.4 Gaussian approximation for the discretized AWT Section 3.2 phase $\arg(W_{Y})$ Section 3.1 magnitude $|W_{Y}|$ Section 3.3 covariance j.p.d.f. (Prop. 1 Cor. 1) contour (Prop. 4) j.p.d.f. (Prop. 5 Cor. 2) stability analysis (Prop. 2-3) stability analysis (Cor. 3) $|W_{Y}|$ (Prop. 6-7 Cor. 4) $\arg(W_{Y})$ (Cor. 5-6) $|W_{Y}|$ vs $\arg(W_{Y})$ (Cor. 7)
@@
The rest of the paper is organized as follows. In Section 2, we present some necessary background, including a review of the continuous wavelet transform and the spectral representation of stationary Gaussian processes. As illustrated in Figure 1, our main findings are presented in Section 3, which is further divided into four subsections addressing: (1) the distributional properties of the magnitude of AWT, (2) the distributional properties of the phase of AWT, (3) the covariance structure of the AWT’s magnitude and phase, and (4) the Gaussian approximation for the discretized AWT of non-Gaussian and non-stationary noise. Section 4 concludes the paper, and Section 5 contains all proofs.
@@ 2. Preliminary
(8) $p\left(\mathbf{w};\mathbf{W}_{Y}\right)=\frac{1}{\pi^{n}\textup{det}(\Gamma)}\textup{exp}\left(-\left(\mathbf{w}-\mathbf{W}_{f}\right)^{*}\Gamma^{-1}\left(\mathbf{w}-\mathbf{W}_{f}\right)\right),$
@@ 3.1. Marginal distributions of AWT magnitude
(18) $p\left(r_{1:n};|\mathbf{W}_{Y}|\right)=\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)\int_{0}^{2\pi}\hskip-4.26773pt\int_{0}^{2\pi}\hskip-5.69046pt\cdots\hskip-4.26773pt\int_{0}^{2\pi}\hskip-2.84544ptp\left(\mathbf{w};\mathbf{W}_{Y}\right)d\theta_{1}d\theta_{2}\cdots d\theta_{n}$
@@
(20) $p\left(r_{1:n};|\mathbf{W}_{Y}|\right)=\frac{1}{\pi^{n}\textup{det}(\Gamma)}\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)J\left(\textup{diag}\left(r_{1:n}\right)\Gamma^{-1}\textup{diag}\left(r_{1:n}\right)\right),$
@@
(22) $p\left(r_{1:2};|\mathbf{W}_{\Phi}|\right)=\frac{4r_{1}r_{2}}{\textup{det}(\Gamma)}\exp\left(-\left(\frac{r_{1}^{2}}{\Gamma_{11}}+\frac{r_{2}^{2}}{\Gamma_{22}}\right)\frac{\Gamma_{11}\Gamma_{22}}{\textup{det}(\Gamma)}\right)I_{0}\left(\frac{2|\Gamma_{12}|r_{1}r_{2}}{\textup{det}(\Gamma)}\right),$
@@
(23) $p\left(r;\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}\right)=2q(t,s)re^{-q(t,s)(1+r^{2})}I_{0}\left(2q(t,s)r\right)$
@@
$p\left(r;\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}\right)=$ $2q(t,s)re^{-q(t,s)(1+r^{2})}\frac{e^{2q(t,s)r}}{\sqrt{4\pi q(t,s)r}}\left(1+\mathcal{O}\left(\frac{1}{2q(t,s)r}\right)\right)$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}-1\right|>\varepsilon\right)\leq\textup{exp}\left(-\frac{\varepsilon^{2}}{16}q(t,s)\right)+\textup{exp}\left(-\frac{\varepsilon}{2}q(t,s)\right).$
@@ 3.2. Marginal distributions of AWT phase
(27) $p\left(\theta_{1:n};\Theta_{Y}\right)=\int_{0}^{\infty}\hskip-4.26773pt\int_{0}^{\infty}\hskip-5.69046pt\cdots\hskip-4.26773pt\int_{0}^{\infty}\left(\overset{n}{\underset{\ell=1}{\prod}}r_{\ell}\right)\hskip-2.84544ptp\left(\mathbf{w};\mathbf{W}_{Y}\right)dr_{1}dr_{2}\cdots dr_{n}$
@@
(28) $p(\theta;\Theta_{Y}(t,s))=\frac{1}{2\pi}e^{-q}+\frac{\sqrt{q}}{2\sqrt{\pi}}\cos(\theta-\theta_{f})\textup{exp}\left\{\hskip-1.42271pt-q\left[1-\cos^{2}(\theta-\theta_{f})\right]\hskip-1.42271pt\right\}\hskip-2.84544ptB_{q}(\theta-\theta_{f}),$
@@
(30) $p\left(\theta_{1:2};\Theta_{\Phi}\right)=\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)}{D(\theta_{2}-\theta_{1})}+\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)C(\theta_{2}-\theta_{1})}{D(\theta_{2}-\theta_{1})^{3/2}}\Big[\frac{\pi}{2}+\tan^{-1}\Big(\frac{C(\theta_{2}-\theta_{1})}{\sqrt{D(\theta_{2}-\theta_{1})}}\Big)\Big],$
@@
$\mathbb{P}\left(d\left(\Theta_{Y}(t,s),\arg(W_{f}(t,s))\right)>\varepsilon\right)\leq\frac{\varepsilon}{\pi}e^{-q}+\frac{2\varepsilon\sqrt{q}}{\sqrt{\pi}}\cos(\varepsilon)\textup{exp}\left\{-q\left[1-\cos^{2}(\varepsilon)\right]\right\},$
@@
$d\left(\Theta_{Y}(t,s),\arg(W_{f}(t,s))\right)=1-\left|\cos\big(\Theta_{Y}(t,s)-\arg(W_{f}(t,s))\big)\right|.$
@@ 3.3. Covariance of AWT’s magnitudes and phases
$=$ $4\underset{\Box,\triangle\in\{\Re,\Im\}}{\sum}\Box(W_{f}(t_{1},s_{1}))\triangle(W_{f}(t_{2},s_{2}))\sqrt{s_{1}s_{2}}\int_{\mathbb{R}}e^{i(t_{1}-t_{2})\lambda}\overline{\widehat{\Box(\psi)}}(s_{1}\lambda)\widehat{\triangle(\psi)}(s_{2}\lambda)F(d\lambda)$
@@
(32) $\textup{Cov}\left(|W_{Y}(t_{1},s_{1})|,|W_{Y}(t_{2},s_{2})|\right)$
$=$ $\int_{\mathbb{C}}|w|p\left(w;W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)dw-\overset{2}{\underset{\ell=1}{\prod}}\int_{\mathbb{C}}|w|p\left(w;W_{Y}(t_{\ell},s_{\ell})\right)dw,$
@@
$p\left(w;W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)$
(37) $=$ $\int_{0}^{\infty}\int_{0}^{2\pi}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\textup{exp}\left(-\left(\left[\begin{array}[]{cc}re^{i\theta}\\ wr^{-1}e^{-i\theta}\end{array}\right]-\left[\begin{array}[]{cc}W_{f}(t_{1},s_{1})\\ W_{f}(t_{2},s_{2})\end{array}\right]\right)^{*}\right$.
@@
(44) $\textup{Corr}\left(\left|W_{\Phi}(t_{1},s_{1})\right|,\left|W_{\Phi}(t_{2},s_{2})\right|\right)\sim\left(\frac{16}{\pi}-4\right)^{-1}\frac{|\Gamma_{12}|^{2}}{\Gamma_{11}\Gamma_{22}}.$
@@
    $\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}=$ $2^{2\alpha+1}\frac{\left(s_{1}s_{2}\right)^{\alpha+\frac{1}{2}}}{(s_{1}+s_{2})^{2\alpha+1}}=\left(\frac{2^{\Delta/2}+2^{-\Delta/2}}{2}\right)^{-(2\alpha+1)}.$
@@
    $\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}=$ $2^{2\alpha+1}\frac{s^{2\alpha+1}}{\left|2s-i(t_{1}-t_{2})\right|^{2\alpha+1}}=\left[1+\left(\frac{t_{1}-t_{2}}{2s}\right)^{2}\right]^{-(\alpha+1/2)}.$
@@
$\mathbb{E}\left[e^{i(\Theta_{\Phi}(t_{1},s_{1})-\Theta_{\Phi}(t_{2},s_{2}))}\right]=\int_{\mathbb{C}}\frac{w}{|w|}p\left(w;W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)dw.$